# Random Scaling of Emergent Capabilities

### v4

@@ Abstract
Language models famously improve under a smooth scaling law, but some specific capabilities exhibit sudden breakthroughs in performance. While advocates of “emergence" view breakthroughs as unlocked capabilities, others attribute them to thresholding effects on noncontinuous metrics. We propose that breakthroughs are instead driven by continuous changes in the probability distribution of training outcomes when performance is bimodally distributed across random seeds. In synthetic length generalization tasks, we show that different random seeds can produce either highly linear or emergent scaling trends. We reveal that sharp breakthroughs in metrics are produced by underlying continuous changes in their distribution across seeds. In a case study of inverse scaling, we show that even as the probability of a successful run declines, the average performance of a successful run increases monotonically. We validate our distributional scaling framework on realistic settings by measuring MMLU performance in LM populations. Our observations hold true even under continuous loss metrics, confirming that random variation must be considered when predicting a model’s performance from its scale.
@@ 1 Introduction
On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales provide the capacity to learn specific concepts (Wei et al., 2022). On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent (Schaeffer et al., 2024). We will argue that such discontinuities are driven by continuous changes in the probability of a breakthrough at each scale. In other words, the discontinuities are real—each model firmly either knows or does not know a given concept—but emergent breakthroughs do not always reflect some fixed threshold which permits the concept to be learned. Instead, models may learn the concept at various scales, albeit with changing probability.
@@
We posit that a breakthrough capability is distinguished not by deterministic responses to scale, but by multimodal random variation. In other words, independent training runs cluster in their performance metrics. This observation is undocumented because scaling laws usually plot a single training run at each scale, rarely testing multiple seeds. Although random variation may be benign when model performance is measured in-distribution (Jordan, 2024), previous work suggests that out-of-distribution performance may vary widely across training runs (Zhou et al., 2024a; b; Qin et al., 2024; Juneja et al., 2023; Li et al., 2025), even at larger scales (Madaan et al., 2024).
@@
By connecting breakthrough scaling with random variation, we challenge the narratives of both the emergence and mirage camps. First, our results contest the position of the mirage or “loss-to-downstream” camp, which claims that effects of scale are predictable and that continuous metrics will smooth out apparently discrete concept learning. We discover clustered multimodal distributions of capabilities, confirming that models unpredictably learn critical discrete concepts—and these concepts are observable even using continuous loss metrics. Furthermore, we complicate the narrative of the competing emergence camp, which often treats specific model sizes as distinct in their capacity. While we confirm that algorithms require some minimum model capacity, that capacity may not be the enormous scale at which emergence is observed. Instead, discontinuous performance jumps are sampled from a continuously changing multimodal distribution, where the “success” component ultimately dominates at larger scales. To extrapolate downstream metrics to large scales, we must predict the likelihood of successful emergence as well as the performance of a successful run.
@@
Figure 1: Different random seeds produce different scaling trends. Scaling trends can be emergent or linear for different seeds, even if all models train on the same data with the same hyperparameters. On the count task, we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right). We mark parameter counts immediately before and after seed 93’s emergence respectively as (a) and (b). Histograms illustrate the bimodal distribution of performance across all random seeds at scales (a) and (b), marking the positions of seeds 93 and 205. Breakthroughs occur when consecutive points represent different clusters; linear trends occur when each point is sampled from the same gradually shifting cluster.
  • Breakthroughs result from bimodal performance distributions. In multiple choice (Section 3.1) and synthetic task length generalization (Section 2.1), some random seeds produce linear scale trends while others are emergent. This variation is caused by the bimodal distribution of each skill across random seeds (Section 2.2), a property that materializes around breakthrough thresholds in model size. At these scales, emergence is a stochastic property (Section 2.3).
  • Bimodal variation persists under continuous metrics. In the emergence debate, one main position (Schaeffer et al., 2024) is that breakthroughs are caused by measuring discontinuous metrics such as accuracy rather than loss. For our tasks, continuous loss metrics can remain visibly bimodally distributed, particularly when there is a more even split of failed and successful runs. We confirm these loss distributions on both synthetic (Section 2.4) and natural (Section 3.4) tasks.
  • When a scale curve exhibits sudden discontinuous improvement in a skill, the probability of learning that skill may be changing continuously. Treating the bimodal distribution as a mix of failure and success distributions, we illustrate that average improvements can come from changes in the probability of success or in the mean performance of a successful run (Section 2.2). Although bimodality can appear abruptly at a minimum capacity scale (Section 2.5), these sudden distributional changes do not necessarily align with breakthrough scales for individual model runs.
  • Competing solutions can lead to either monotonic or U-shaped trends in emergence likelihood. We discuss a synthetic task that exhibits periods of inverse scaling in average improvement (Section 2.6), a phenomenon also observed in LMs (McKenzie et al., 2022). We note that the mean performance of successful runs remain monotonic during ranges of inverse scaling, suggesting that U-shaped curves in performance distributions are an artifact of success probability.
  • In realistic settings, multimodal performance clusters emerge based on data composition and model scale. In Section 3, we extend our analysis to the task of multiple choice question answering by training partially reinitialized LMs. Our findings confirm that random variation in LMs is also bimodal for established emergent capabilities near their breakthrough threshold.
@@
2 Synthetic Length Generalization Experiments
Usually, performance breakthroughs are measured on a single model per scale or, at most, the average of a few runs. The literature suggests that these emergent capabilities unlock at specific model scales (Wei et al., 2022), implying that scaling curves for different seeds would perform similarly to within a margin of noise. Contrary to this belief, we demonstrate that performance is sampled from a stochastic distribution which changes gradually even as individual scaling curves jump abruptly.
@@ 2.1 Tasks and Setup
After training models on synthetic tasks, we measure their length generalization (Graves et al., 2016; Kaiser and Sutskever, 2015; Lake and Baroni, 2018; Hupkes et al., 2020), one of many compositional properties that can lead to conceptual breakthroughs (Srivastava et al., 2023; Löwe et al., 2024; Chen et al., 2024). Further experimental details are in Appendix A.
@@
Task: We consider two algorithmic tasks previously studied in Zhou et al. (2024a): counting and reverse order addition. Performance on these task ranges widely across random seeds. At each scale, we train models from 250 seeds for the count task and 200 seeds for reverse order addition.
  • Count: Given two numbers in increasing order, the model is trained to generate a sequence which counts consecutively from the first number to the second number. Examples are given in the form "5, 9 >, 5, 6, 7, 8, 9", while limiting the length of the counting sequence during training. Zhou et al. (2024a) showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
  • Addition: Zhou et al. (2024b) showed that Transformers can generalize 10-15 digits past training length for an addition task, if provided with index hints and allowed to generate the answer backwards. Examples take the form "a0, 3, a1, 4, +, a0, 2, a1, 8, >, a1, 2, a0, 6", where index hints are a consecutive sequence of the max number of digits sampled randomly from 0 to the max evaluation length. We use this modified form of reverse order addition with index hints as our addition task.
@@ 2.2 Emergence is a Sign of Bimodal Variation
What do the scaling curves for length generalization tasks look like when we generate them in conventional ways? Following Srivastava et al. (2023), we take the vector of model performances for the count task at test length 60 across different scales with a fixed initialization and shuffle seed. [^1] Although it is standard practice to fix the random seed when reporting LM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales. We calculate the task’s breakthroughness and linearity metrics on each seed. As defined in Appendix B, the former metric measures emergence, whereas the latter measures a smooth response to scale. We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Figure 1. As seen in Figure 1 and Appendix Figure 10, we can easily find fixed seeds that show varying levels of emergence and linearity, due to random variation in breakthroughs.
This variation is explained by the bimodality of model performance distributions when varying seed. Figure 2 illustrates that, for a population of models independently trained on the addition task, length generalization clusters into high and low component modes at many parameter sizes. This clustering produces distinctly bimodal performance distributions, causing some model runs to appear as breakthroughs while others follow a more linear progression. When bimodally distributed runs cluster into very high and very low performance components, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster. These differences ultimately lead to high variability in the timing and degree of emergence. Furthermore, these differences can be a source of oscillating scale curves, as Srivastava et al. (2023) documented in their lowest-linearity tasks.
Figure 3: Summary statistics for length generalization (addition task). Exact match statistics for 200 models trained at length 35 and tested at length 40. We track overall 3 mode and 3 mean. Because the EM accuracy distribution is bimodal, the mode exhibits a sharp increase even as the mean evolves continuously. Defining success as >20% EM accuracy, we also note the continuous change in 3 the probability of success and in 3 the mean of successful runs. Means feature 95% confidence intervals over 1000 bootstrapped samples.
@@ 2.3 Sudden Jumps From Gradual Distribution Shifts
When reporting metrics from only one seed or the mean of a few seeds, the outcome is likely to be close to the mode performance of the underlying model population. As shown in Figure 3 (top left), the mode of the performance distribution shows a massive spike in improvement mirroring an emergent benchmark’s scaling trend. However, we claim that this discontinuous improvement is only an artifact of sampling the mode performance, although the underlying bimodality also produces discontinuity in many single-seed scaling curves. Underlying this discontinuous performance jump are continuous changes in other distributional statistics. In Figure 3 (top right), the mean exhibits a smoother trend in accuracy. Mode and mean diverge thus because the underlying distribution is bimodal, expressing a mixture of “successful” and “failing” runs.
Treating the distribution as a mixture of successes and failures, we can separately analyze the probability of a successful run and the performance distribution of successful runs. Both of these properties are changing continuously and gradually when the mode increases abruptly. If we restrict our analysis to the runs achieving nontrivial 20% accuracy, we see that the probability (Figure 3 (bottom left) and mean (bottom right) of such “successful” runs both exhibit continuous improvement, with the exception of increasing from depth 2 to 3 (discussed further in Section 2.5). Even at the mode breakthrough,these underlying distributional properties are only changing gradually. We conclude that when tasks exhibit bimodal distributions across random seeds, there are statistics exhibiting continuous improvements underlying the seemingly abrupt improvements across scale.
@@ 2.4 Is Bimodality a Mirage?
Figure 4: Random variation in continuous length generalization error (addition task). Kernel Density Estimation (KDE) of our loss-based error metric (Equation 1) across model runs. At scales where the EM accuracy distribution is bimodal, the distribution remains bimodal even when using a continuous metric.
Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior (Schaeffer et al., 2024); conversely, continuous metrics turn apparent emergence into smooth curves (Srivastava et al., 2023). We must be particularly cautious about claiming emergence when requiring outputs to exactly match a target string, as we have so far. Are our case studies artifacts of thresholding effects? Do our bimodal distributions become unimodal under continuous metrics?
To avoid thresholding artifacts, we use a continuous equivalent to the exact match metric: the maximum loss assigned to any individual token. Each token is individually computed in the correct output context, so the entire sequence represents a fixed set of continuous per-token loss functions. Because the maximum of a fixed set of Lipschitz-continuous functions is always Lipschitz-continuous, this error score is guaranteed to be continuous as long as per-token loss is continuous. For model $f$ on dataset $X$, the continuous error score based on per-token loss $L(f(x_{0\ldots i-1}),x_{i})$ is:
In Figure 4, we plot this continuous metric for addition models on the length generalization test dataset. The distribution of this metric across random seeds is often still clustered. We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric; their bimodality is not due to thresholding alone. Further discussion around continuous metric plots for count and histograms for probability-based metrics are in Appendix D.
2.5 Bimodality Emerges Abruptly
(a) Addition task. The sharpest change in distance occurs early. During the mode breakthrough, the distributions are still shifting gradually.
(b) Count task. The distribution changes slowly between intermediate model scales, but changes suddenly at the nadir of the U-shaped curve in Figure 6.
Figure 5: Changes in random variation. Wasserstein-L2 distance of each scale’s performance distribution relative to the largest scale, scaling depth and width independently. We mark the emergence of bimodality at the last scale before multiple peaks appear. We mark the mode breakthrough at the last scale before successful length generalization becomes marginally more likely than failure.
@@ 2.5 Bimodality Emerges Abruptly
The previous section showed that bimodal performance distributions produce emergent scaling curves for individual seeds. We next analyze how the distribution itself evolves across scales. As shown in Figure 2, the performance distribution starts as unimodal at the smallest scale, where no models can length generalize. Larger scales yield a bimodal distribution where most of the probability mass is ultimately placed on successful length generalizing runs. A priori, there could be a smooth evolution between these two distributions in which probability mass from failing runs gradually shifts towards higher performance metrics, eventually splitting into a clear separate cluster. We find instead that the shift from unimodal to bimodal is abrupt and instantly polarized into low and high clusters. In Appendix Figure 12 we show that models fail to generalize when fixing depth to be one layer even when scaling width (despite having perfect in-distribution accuracy).
@@
To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each distribution relative to that of the final, largest model scale when separately fixing width and depth. In Figure 5(a) we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth reaches 3 layers or model width reaches 2 heads. These sudden changes identify the moment when a new capability is unlocked, as the distribution transitions abruptly from unimodal on one extreme to bimodal at both extremes. We posit that this transition marks the minimum capacity required to learn the task. We also mark the point in each trend where the mode in Figure 3 (top left) increases sharply. We posit that it may be misleading to draw conclusions about minimal model capacity on a specific task using single runs at each scale, whereas distributional metrics correctly predict the sudden appearance of probability mass placed on successful runs.
2.6 Count: An Instance of U-shaped Scaling
We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Appendix Figure 11) but yields a very different scaling effect: a U-shaped curve. Figure 6 reveals this peculiar phenomenon in the mean accuracy scaling curve, when holding depth fixed. This curve is not simply a result of the summary statistic chosen, as it is mirrored by the evolution of the distributions as a whole according to their W2 distance in Figure 5(b).
U-shaped scaling has been observed in LMs, but its causes are not currently well-understood. When the Inverse Scaling Prize (McKenzie et al., 2022) solicited tasks which exhibit inverse scaling trends—performance decreasing with scale—for large models, Wei et al. (2023) revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models. Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution. Indeed, Figure 6 (bottom right) shows that although the trend in the mean across all runs is U-shaped curve, the mean of the “successful” runs—those achieving at least 50% accuracy—still improves monotonically when increasing width. The observation of inverse scaling is, instead, due to changes in the probability of success (bottom left). Even when inverse scaling is in effect across a performance distribution, the performance of successful runs may exhibit more conventional responses to scale.
Figure 6: Summary statistics for length generalization (count task). Models tested at length 60, scaling width (with fixed depth 4) and depth (with fixed width 4) independently. The trend for fixed depth (scaling width) exhibits a U-shaped curve for the mean across all runs 6; however, the mean of successful runs (defined as >50% EM accuracy) 6 still improves continuously.
(a) Qwen2.5-0.5B trained on different mixes.
(b) Qwen2.5 models trained on 10% MMLU data.
Figure 7: Performance on MMLU. All models have their last layer reinitialized randomly and then are trained on a mix of C4 news and MMLU, with varying proportions of MMLU. Left: With insufficient MMLU training data, each model fails to follow the multiple-choice format, resulting in trivial performance. As more MMLU data is included, different seeds lead to bimodally distributed performance. With enough MMLU data, all models score above the random baseline. Right: For a fixed data mix, the smaller model has bimodally distributed performance. In contrast, the larger model consistently relearns to answer multiple-choice questions, though not at its original base performance.
3 LLM Experiments
After connecting emergence with bimodality in small synthetic settings, we turn to LMs. We focus on the MMLU dataset, where high performance emerges after the LM learns the multiple choice format (Srivastava et al., 2023; Hu and Frank, 2024). To avoid the expense of repeatedly training large-scale models from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LMs before continuing to train them. Because Qwen base models perform well for their scale (39.5% for Qwen2.5-0.5B), these experiments effectively require models to re-discover MMLU capabilities after their removal. Our continued training dataset, a mix of C4 news data and the MMLU auxiliary training data, is insufficient to recover the full MMLU capability of the the base model.

### v5

@@ Abstract
Language models famously improve under a smooth scaling law, but some specific capabilities exhibit sudden breakthroughs in performance. Advocates of “emergence” view these capabilities as unlocked at a specific scale, but others attribute breakthroughs to superficial metric thresholding effects. We propose that breakthroughs are instead driven by continuous changes in the probability distribution of training outcomes when performance is bimodally distributed across random seeds. we show that different random seeds can produce either smooth or emergent scaling trends in synthetic length generalization tasks, multiple choice question answering, and grammatical generalization. We reveal that sharp breakthroughs in metrics are produced by underlying continuous changes in their distribution across seeds. These distributions may become abruptly bimodal at a capacity threshold—but this threshold appears at scales well before most seeds achieve breakthrough. Our observations hold true even under continuous loss metrics, confirming that random variation must be considered when predicting a model’s performance from its scale.
Machine Learning, ICML
@@ 1 Introduction
On one side, advocates of emergence claim that performance abruptly improves when a particular scale provides the capacity to learn specific concepts (Wei et al., 2022). On the other side, skeptics argue that these sudden improvements are a mirage (Schaeffer et al., 2024) driven by thresholding effects. These threshold artifacts are alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent. We argue that discontinuities are driven by continuous changes in the probability of a breakthrough at each scale. In other words, the discontinuities are real—each model firmly either knows or does not know a given concept—but breakthroughs do not always reflect a fixed threshold at which a concept is learnable. Instead, models may learn the concept at various scales, albeit with changing probability.
Figure 1: Different random seeds produce different scaling trends. Scaling trends can be emergent or linear for different seeds, even if all models train on the same data with the same hyperparameters. On the count task (see Appx. C), we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right). We mark parameter counts immediately before and after seed 93’s emergence, respectively, as (a) and (b). Histograms illustrate the bimodal distribution of performance across all random seeds at scales (a) and (b), marking the positions of seeds 93 and 205. Breakthroughs occur when consecutive points represent different clusters; linear trends occur when each point is sampled from the same gradually shifting cluster.
@@
We posit that a breakthrough capability is distinguished not by deterministic responses to scale, but by multimodal random variation. In other words, independent training runs cluster in their performance metrics. This observation is undocumented because scaling laws usually plot a single training run at each scale, rarely testing multiple seeds. Although random variation may be benign when model performance is measured in-distribution (Jordan, 2024), previous work suggests that out-of-distribution performance may vary widely across training runs (Zhou et al., 2024a, b; Qin et al., 2024; Juneja et al., 2023; Li et al., 2025), even at larger scales (Madaan et al., 2024).
@@
By connecting breakthrough scaling with random variation, we challenge the narratives of both the emergence and mirage camps. First, our results contest the position of the mirage or “loss-to-downstream” camp: that effects of scale are predictable and that continuous metrics will smooth out apparent breakthroughs. We discover clustered multimodal distributions of capabilities, confirming that models unpredictably learn critical discrete concepts—and these clusters are observable even using continuous loss metrics. Furthermore, we complicate the rival narrative of the emergence camp, which often treats specific model sizes as distinct in their capacity. While we confirm that algorithms require some minimum model capacity, that capacity may not be the enormous scale at which emergence is observed. Instead, discontinuous performance jumps are sampled from a continuously changing multimodal distribution, where the “success” mode ultimately dominates at larger scales. To extrapolate downstream metrics to large scales, we must predict the likelihood of successful emergence as well as the performance of a successful run.
@@
  • Breakthroughs result from bimodal performance distributions. On synthetic algorithmic tasks (Sec. 2), simple language modeling tasks (Sec. 4) and multiple choice question answering (Sec. 3) , some random seeds produce linear scale trends while others are emergent. This variation is caused by the bimodal distribution of each skill across seeds (Sec. 2.2), a property that materializes around breakthrough thresholds in model size. At these scales, emergence is a stochastic property (Sec. 2.3).
  • Bimodal variation persists under continuous metrics. In the emergence debate, one main position (Schaeffer et al., 2024) is that breakthroughs are caused by measuring discontinuous metrics such as exact match accuracy rather than loss. We find continuous loss metrics can remain visibly bimodally distributed, particularly when there is a more even split of failed and successful runs. We confirm these results on both synthetic (Sec. 2.4, Sec. 4.2) and natural (Sec. 3.3) tasks.
  • When a scale curve exhibits sudden discontinuous improvement in a skill, the probability of learning that skill may be changing continuously. Treating the bimodal distribution as a mix of failure and success distributions, we illustrate that average improvements can come from changes in the probability of success or in the mean performance of a successful run (Sec. 2.2). Although bimodality can appear abruptly at a minimum capacity scale (Sec. 2.5), these sudden distributional changes do not necessarily align with breakthrough scales for individual model runs.
  • Bimodality can produce nonmonotonic scaling trends. We discuss a synthetic task with ranges of inverse scaling in mean accuracy (Appx. C.2). The average of successful runs remains monotonic in these ranges, suggesting that U-shaped curves in overall mean are an artifact of success probability.
2 Synthetic Length Generalization Tasks
Usually, scaling trends are measured on a single model per scale or, at most, the average of a few runs. The literature suggests that emergent capabilities unlock at specific model scales (Wei et al., 2022), implying that different training runs would perform similarly at each scale. Contrary to this belief, we demonstrate that performance is sampled from a stochastic distribution which changes gradually even as individual scaling curves jump abruptly.
@@ 2.1 Tasks and Setup
After training models on synthetic tasks, we measure their length generalization (Graves et al., 2016; Kaiser and Sutskever, 2015; Lake and Baroni, 2018; Hupkes et al., 2020), one of many compositional skills that can lead to conceptual breakthroughs (Srivastava et al., 2023; löwe2024abruptspontaneousstrategyswitches; Chen et al., 2024). Experimental details are in Appx. A.
@@
Task: We consider two algorithmic tasks previously studied in Zhou et al. (2024a): counting and addition. An analysis of the count task is provided in Appx. C, including discussion (Appx. C.2) of how its average accuracy responds non-monotonically to scale. In the main body of this paper, we will focus on the addition task. Zhou et al. (2024b) showed that Transformers can generalize 10-15 digits past training length for an addition task, if provided with index hints and allowed to generate the answer backwards. We use this modified reverse-order addition task.
@@ 2.2 Emergence is a Sign of Bimodal Variation
What do the scaling curves for length generalization tasks look like when we generate them in conventional ways? Following Srivastava et al. (2023), we calculate the breakthroughness and linearity of scaling curves where each run shares a fixed initialization and shuffle seed. [^1] Although it is standard practice to fix the random seed when comparing LM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation between different model sizes. As defined in Appx. B, breakthroughness measures emergence, whereas linearity measures a smooth response to scale. We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Fig. 1 (with other extremes shown in Appx. Fig. 10). Because breakthroughs vary across seeds, we can easily find fixed seeds that lead to either emergent or smooth scaling.
This variation is explained by the bimodality of model performance distributions when varying seed. Fig. 2 illustrates that, for a population of models independently trained on the addition task, length generalization ability clusters into high and low component modes at many parameter sizes. This clustering produces distinctly bimodal performance distributions, causing some model runs to appear as breakthroughs while others generalize poorly. When bimodally distributed runs cluster into distinct high- and low-performance components, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster. These differences ultimately lead to high variability in the timing and degree of emergence. Furthermore, these differences cause oscillating scale curves like those seen in low-linearity tasks (Srivastava et al., 2023, ref. Fig. 7c).
Figure 3: Summary statistics for length generalization accuracy on the reverse-order addition task. We track overall mode (top left) and overall mean (top right). While the mode exhibits a sharp increase in accuracy at a certain scale of width or depth, the mean evolves more continuously as a result of the bimodal nature of the random variation distribution. We also study successful runs, plotting the fraction of runs reaching above 20% accuracy (bottom left) and the mean of such runs (bottom right). The plots for mean include 95% confidence intervals with 1000 bootstrapped samples.
@@ 2.3 Sudden Jumps From Gradual Distribution Shifts
When reporting metrics from only one seed or the mean of a few seeds, the outcome is likely to be close to the mode performance of the underlying model population. In Fig. 3 (top left), the mode performance shows a massive spike, mirroring an emergent benchmark’s scaling trend. However, we claim that this discontinuous improvement in mode—as well as the discontinuities seen in single-seed emergent scaling curves—is an artifact of bimodality. Underlying this discontinuous performance jump are continuous changes in other distributional statistics. In Fig. 3 (top right), the mean exhibits a smoother trend in accuracy. Mode and mean differ because the underlying distribution is bimodal, expressing a mixture of “successful” and “failing” runs.
Treating the distribution as a mixture of successes and failures, we can separately analyze the probability of a successful run and the performance distribution of successful runs. Both of these properties are changing continuously and gradually when the mode increases abruptly. If we restrict our analysis to the runs achieving at least nontrivial 20% accuracy, we see that the probability (Fig. 3 (bottom left)) and mean (bottom right) of such “successful” runs both exhibit continuous improvement, with the exception of increasing from depth 2 to 3 (discussed further in Sec. 2.5). Even at the mode breakthrough, these underlying distributional properties are only changing gradually. We conclude that when tasks exhibit bimodal distributions across random seeds, gradual statistical improvements cause seemingly abrupt improvements with scale.
@@ 2.4 Is Bimodality a Mirage?
Metrics with hard thresholds can artificially induce breakthroughs (Schaeffer et al., 2024); conversely, continuous metrics turn apparent emergence into smooth curves (Srivastava et al., 2023). We must be particularly cautious about claiming emergence when requiring outputs to exactly match a target string, as we have so far. Are our case studies artifacts of thresholding effects? Do our bimodal distributions become unimodal under continuous metrics?
To avoid thresholding artifacts, consider a continuous equivalent to the exact match metric: the maximum loss assigned to any individual token. Each token is individually computed in the correct output context, so the entire sequence represents a fixed set of continuous per-token loss functions. Because the maximum of a fixed set of Lipschitz-continuous functions is always Lipschitz-continuous, this error score is guaranteed to be continuous as long as per-token loss is continuous. For model $f$ on dataset $X$, the continuous error score based on per-token loss $L(f(x_{0\ldots i-1}),x_{i})$ is:
In Fig. 4 (a)(b), we plot this continuous metric for addition models on the length generalization dataset. The distribution of this metric across random seeds is still clustered. We therefore confirm that emergent capabilities exhibit bimodal distributions even when using a continuous performance metric; their bimodality is not due to thresholding alone. Further findings from these experiments are in Appx. D.
2.5 Bimodality Emerges Abruptly
@@
(c) Changes in random variation.
Figure 4: Analysis of performance distributions (addition task). Left (a, b): Kernel Density Estimation (KDE) of our loss-based error metric (Eq. 1). At scales where accuracy is bimodal, the distribution remains bimodal even with a continuous metric. Right (c): Wasserstein-L2 distance of each scale’s performance relative to the largest scale. We mark the “emergence of bimodality” and the “mode breakthrough.” During breakthrough, distributions shift gradually.
@@
The previous section showed that bimodal performance distributions produce emergent scaling curves for individual seeds. We next show that the distribution itself can change suddenly at certain scales. As shown in Fig. 2, the performance distribution starts as unimodal at the smallest scale, where no models can length generalize. Larger scales yield a bimodal distribution where most of the probability mass is ultimately placed on successful (i.e., length-generalizing) runs. A priori, there could be a smooth evolution between these two distributions in which probability mass from failing runs gradually shifts towards higher performance metrics, eventually splitting into a clear separate cluster. We find instead that the shift from unimodal to bimodal is abrupt and instantly polarized into low and high clusters. In Appx. Fig. 12, we confirm that models fail to generalize when fixing depth to be one layer even when scaling width (despite having perfect in-distribution accuracy), illustrating that the length-general function is suddenly learnable at sufficient depth.
@@
To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each scale’s distribution relative to that of the final, largest model scale. In Fig. 4c we see a sharp decrease in the W2 distance, marking the sudden appearance of highly successful runs when model depth reaches 3 layers or model width reaches 2 heads. These sudden changes identify the moment when a new capability is unlocked, as the distribution transitions abruptly from unimodal to bimodal. We posit that this transition marks the minimum capacity required to learn the task. We also mark the point in each trend where the mode in Fig. 3 (top left) increases sharply. We argue that one can draw misleading conclusions about minimal model capacity when only studying using single runs at each scale, whereas distributional statistics correctly identify the minimum capacity that a skill requires.
3 Large Language Model Experiments
After connecting emergence with bimodality in small synthetic settings, we turn to LMs. We focus on the MMLU multiple-choice question answering dataset, where performance jumps (Srivastava et al., 2023) after the LM learns the multiple choice format (Hu and Frank, 2024). We will show that emergent scaling curves in LMs express underlying multimodal performance distributions.